OpenStudy (anonymous):

right triangle using pythagorean theorem: a^2+(x+2)^2=(2x-2)^2

OpenStudy (watchmath):

what is the instruction?

OpenStudy (anonymous):

find 'a' using pythagorean theorem

OpenStudy (anonymous):

a=3(x^2 -4x)

OpenStudy (anonymous):

a^2 + b^2 = c^2 so (2x-2)^2-(x+2)^2= a^2

OpenStudy (anonymous):

hmm i tried that answer and it told me it was wrong..

OpenStudy (anonymous):

did it ask you to find a or a^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it asked you to find a not a^2

OpenStudy (anonymous):

yes, at the end you get rid of the ^2

OpenStudy (anonymous):

if u square root the answer i gave you you will end up with the right one

OpenStudy (anonymous):

is the answer -x? after you square root it?